The denominator of a fraction is 4 less than the numerator. If the denominator is doubled and the numerator is increased by 6, the sum of the original fraction and the new one is 3. Find the original fraction. Thank you
Answer by mananth(16946)

The denominator of a fraction is 4 less than the numerator. If the denominator is doubled and the numerator is increased by 6, the sum of the original
fraction and onethe new is 3. Find the original fraction. Thank you
Let the numerator be x
the denominator will be x-4
The numerator is increased by 6 so numerator will be x+6
The denominator will be 2(x-4)
x / x-4 + x+6/2(x-4) =3
2x +x+6 /2(x-4) =3
3x+6 =6(x-4)
3x+6 =6x-24
3x-6x=-24-6
-3x=-30
x=10 which is the numerator
denominator will be 10-4=6
the fraction is 10/6